OpenStudy (anonymous):

Kathryn, a college student, works part time at a coffee house. She worked 3 3/20 hours on monday, Tuesday 6 1/2, Wed 2 3/4, Thurs 5 1/4 Friday 8 hours. How many hours did she work all together.

OpenStudy (johnweldon1993):

Well lets see \[\frac{63}{20} + \frac{13}{2} + \frac{11}{4} + \frac{21}{4} + 8\] The least common denominator = 20 so \[\frac{63}{20} + \frac{130}{20} + \frac{55}{20} + \frac{105}{20} + \frac{160}{20}\] Add them \[\large \frac{63 + 130+55+105+160}{20}\] \[\large \frac{513}{20}\] Now just divide it by 20 to change this back into a mixed number \[\large 25 \frac{13}{20}\]

OpenStudy (anonymous):

TY